OpenStudy (anonymous):

Use the following equation: log2x + log2(x – 7) = 3 The answer to the function would be: Select one: a. 8 b. –1 c. 8 or –1 d. –8 or 1

OpenStudy (anonymous):

You can use here: \[\log(a) + \log(b) = \log(a \times b)\]

OpenStudy (campbell_st):

well here is a tip... you can't find the log of a negative number.... so that should help identify the answer

OpenStudy (princeharryyy):

is 2 in base @chantz417

OpenStudy (anonymous):

Yes sorry just got back

OpenStudy (princeharryyy):

where 2 is in base of logarithm

OpenStudy (princeharryyy):

u there @chantz417

OpenStudy (campbell_st):

lol... how can you substitute x = -1 what is the value of \[\log_{2}(-1) =..?\]

OpenStudy (campbell_st):

wow... understand some basics of logarithms

OpenStudy (princeharryyy):

you can after u make it log(x(x-7)) after that u get x(x-7) > 0 implying x can be either greater then 7 or less than -1 read logarithms and functions carefully @campbell_st

OpenStudy (princeharryyy):

u can just put x=-1 in logx that's wrong but you can after u make x(x-7) as it will come out positive @campbell_st

OpenStudy (princeharryyy):

can't* not can typed wrongly

OpenStudy (campbell_st):

have a look at the original equation you get \[\log_{2}( -1) + \log(-1 - 7) = 3\] tell me how that works...

OpenStudy (campbell_st):

the only solution is 8....

OpenStudy (princeharryyy):

seperately it doesn't but if u make it something like this log(-1)(-1-7)

OpenStudy (princeharryyy):

so i missed u clicked u are right we must see it other perspective as well @campbell_st gud luck !!!

OpenStudy (campbell_st):

just test the solution in the original equation... the only solution that works is x = 8

OpenStudy (princeharryyy):

ans 8 only

OpenStudy (campbell_st):

you are provided with the answer choices.... the easy solution with out calculation or algebraic manipulation is substitution... so substitute the choices and see what you get

OpenStudy (princeharryyy):

not every one is that intelligent @campbell_st you need to xplain others

OpenStudy (princeharryyy):

u and i can do that but not every one

OpenStudy (campbell_st):

well I'm in high school... and have been taught to always check the solutions make sense.

OpenStudy (princeharryyy):

it's ohk but doesn't always work that way @campbell_st still u have a long way to go

OpenStudy (princeharryyy):

u will get options sometimes u can't just check them out

OpenStudy (campbell_st):

lol... why make things hard.... use the obvious... another good problem solving skill.

OpenStudy (princeharryyy):

u'll learn why to solve ... later it might look that easy today .... @campbell_st